OpenStudy (anonymous):

what is the fourth tem in this sequence? tn=3.2n-5 a.-7.2 b.-16.8 c.7.8

OpenStudy (anonymous):

You ask a lot of questions lol. Assuming n starts at 1... t(4)=3.2(4)-5=12.8-5=7.8

OpenStudy (anonymous):

i know lol this is a big assignment i need help with or i fail my algebra class